OpenStudy (anonymous):

If B is between A and C find the value of BC. AB = 3, BC = 5x - 1, AC = 12

OpenStudy (aroub):

ab+bc=ac 3+5x=12 5x=-3+12 5x=9 x=9/5

OpenStudy (aroub):

right?

OpenStudy (anonymous):

I guess small correction @ aroub 3+5x-1=12 5x=10 x=2

OpenStudy (anonymous):

I mean BC=2

OpenStudy (aroub):

oh! ya ur correct sorry my mistake

OpenStudy (aroub):

small correction@ lakshmandonala bc=10